Hi, I am trying to come up with a equation for this word problem
Ten years from now Alex's age will be 3 times her present age. Find Alex's age now.
I let (x) be Alex's age now, and I figured her age 10 yrs from now would be 3(x+10) therefore I said x=3(x+10)
Is this correct? Please help me. Thanks

Answer by stanbon(75887)

Ten years from now Alex's age will be 3 times her present age. Find Alex's age now.
Equation:
Let Alex's age now be "A":
(A+10) = 3A
2A = 10
A = 5 yrs. old now
Cheers,
Stan H.
------
I let (x) be Alex's age now, and I figured her age 10 yrs from now would be 3(x+10) therefore I said x=3(x+10)
Is this correct?
-----
Comment: 3(x+10) is "3 times her age 10 years from now"
You want 3 times her age now.